OpenStudy (anonymous):

Vector A has a magnitude of 11 units, vector B has a magnitude of 8.8 units, and A · B has a value of 14.0. What is the angle between the directions of A and B? Vector A has a magnitude of 11 units, vector B has a magnitude of 8.8 units, and A · B has a value of 14.0. What is the angle between the directions of A and B? @Physics

OpenStudy (anonymous):

the dot product of a and b can be rewritten as a*b * cos(x) 14.0 = 11 * 8.8 * cos(x) 14 / 96.8 = cos(x) cos^-1 (.144) = x 81.68 = x

OpenStudy (anonymous):

note that the same process can be applied to a cross product, however the trig function is then a sin